In a youngs double slit experiment the ratio of light mxima to minina close to the central fringe on the screen is 9:1. The ratio of slit width is
1
Expert's answer
2020-05-04T12:40:16-0400
"\frac{9}{1}=\frac{I_{max}}{I_{min}}=\frac{(\sqrt{I_{2}}+\sqrt{I_{1}})^2}{(\sqrt{I_{2}}-\sqrt{I_{1}})^2}"

"\frac{3}{1}=\frac{\sqrt{\frac{I_2}{I_1}}+1}{\sqrt{\frac{I_2}{I_1}}-1}"

"\sqrt{\frac{I_2}{I_1}}=2"

"\frac{I_2}{I_1}=4"